A collection of triangles and pentagons contains 40 shapes having a total of 138 sides. How many pentagons are there?
Answer by stanbon(75887)

A collection of triangles and pentagons contains 40 shapes having a total of 138 sides. How many pentagons are there?
----------
# of shapes equation: p + t = 40
# of sides equation: 5p +3t = 138
--------------------------
Multiply thru 1st by 3 to get:
3p + 3t = 120
5p + 3t = 138
----------------
Subtract 1st from 2nd to get:
2p = 18
p = 9 (# of pentagons)
Since p + t = 40, t = 31 (# of triangles)
=====================
Cheers,
Stan H.